Tim Hudson of the Oakland Athletics had an ERA of 3.53 in 2004, At that Rate, how many runs would he be expected to allow in 100 innings pitched? Round your answer to the nearest whole number.
Answer by checkley77(12844)

The average of 3.53 ERA is 3.53 runs per 9 innings pitched.
100/9=11.11 total games.
11.11*3.53=39.2 runs for 100 innings or 11.11 games pitched.
----------------------------------------
100/9*3.53=353/9=39.2 runs are expected.